# Subluminal and superluminal velocities of free-space photons

### v1

@@ Abstract
We consider rectilinear free-space propagation of electromagnetic wavepackets using electromagnetic field theory, scalar wavepacket propagation, and quantum-mechanical formalism. We demonstrate that spatially localized wavepackets are inherently characterized by a subluminal group velocity and a superluminal phase velocity, whose product equals $c^{2}$. These velocities are also known as the ‘energy’ and ‘momentum’ velocities, introduced by K. Milton and J. Schwinger. We illustrate general conclusions by explicit calculations for Gaussian beams and wavepackets, and also highlight subtleties of the quantum-mechanical description based on the ‘photon wavefunction’.
@@ I Introduction
Subluminal and superluminal velocities in the propagation of light have intrigued scientists for several decades Brillouin (1960). Multiple distinct velocities characterize light propagation, such as phase, group, signal, and energy-transfer velocities. Furthermore, various physical mechanisms can lead to subluminal or superluminal behavior: interactions with matter Chiao and Steinberg (1997); Winful (2006); Boyd and Gauthier (2002); Khurgin (2010); Asano et al. (2016), nontrivial quantum-vacuum effects Latorre, Pascual, and Tarrach (1995), local phase gradients in structured free-space light Berry (2010, 2012); Bliokh et al. (2013), specially engineered space-time wavepackets Bliokh and Nori (2012a); Kondakci and Abouraddy (2019); Yessenov et al. (2022), and the tilt of plane-wave Fourier components in transversely confined light beams Horváth et al. (1996); Porras (2002); Porras, Gonzalo, and Mondello (2003); Giovannini et al. (2015); Bouchard et al. (2016); Gouesbet and Lock (2016); Bareza and Hermosa (2016); Saari (2018); Petrov (2019); Bliokh (2023a).
In this work, we address the simplest scenario among these, namely, the free-space propagation of electromagnetic wavepackets without engineered internal structure. The only requirement is suitable spatial confinement of the wavepacket, which is necessarily entails a corresponding spread in momentum (wavevector) space. As theoretically described and experimentally observed in Giovannini et al. (2015); Bouchard et al. (2016); Bareza and Hermosa (2016); Bliokh (2023a), such wavepackets or single photons propagate with a subluminal group velocity $v_{g}<c$, determined by the transverse confinement of the wavepacket. In addition, we show that the wavepacket can also be characterized by a superluminal phase velocity $v_{ph}>c$, such that $v_{g}v_{ph}=c^{2}$, see Fig. 1.
We trace the origin of this result through several complementary approaches and establish useful connections between relativistic field theory, quantum mechanics, and previous studies of this phenomenon.
@@ III Classical wavepacket considerations
This relation is fully analogous to Eq. (4) for the ‘momentum velocity’, so that $\langle{\bf v}_{ph}\rangle\cdot\langle{\bf v}_{g}\rangle=c^{2}$.
@@
where $z_{R}=kw_{0}^{2}/2$ is the Rayleigh range, i.e., the characteristic longitudinal scale of beam diffraction. Equation (8) means that a paraxial Gaussian wavepacket accumulates a half-wavelength retardation after propagating a distance $2\pi z_{R}$, i.e., about six Rayleigh ranges, see Fig. 2. This estimate agrees with Refs. Giovannini et al. (2015); Bareza and Hermosa (2016), but the retardation effect calculated in Ref. Bliokh (2023a) is twice as small. This difference arises because Ref. Bliokh (2023a) considered a 2D Gaussian wavepacket in the $(x,z)$ plane, whereas here we treat a fully 3D beam. The number of transverse dimensions is doubled, and so is the resulting correction.
@@
$\psi(r_{\perp},z,t)\propto\frac{w_{0}}{w(z)}\exp\!\left({-r_{\perp}^{2}}/w(z)^{2}\right)\times$
$\exp\!\left(ikz+i\frac{kr_{\perp}^{2}z}{2(z_{R}^{2}+z^{2})}-i\arctan(z/z_{R})-i\omega t\right)\!.$ (10)
@@ IV Quantum-mechanical formalism
where the bra-ket inner product implies spatial integration, and we used ${\bf F}^{*}\cdot{\bf F}\equiv|{\bf F}|^{2}=({\bf E}^{2}+{\bf H}^{2})/2=U$. Thus, Eq. (15) is fully equivalent to Eq. (2) for the energy-centroid velocity ${\bf v}_{E}$.
@@ V Peculiarities related to the photon wavefunction
VI Conclusions

### v2

@@ Abstract
We consider rectilinear free-space propagation of electromagnetic wavepackets using electromagnetic field theory, scalar wavepacket propagation, and quantum-mechanical formalism. We demonstrate that spatially localized wavepackets are inherently characterized by a subluminal group velocity and a superluminal phase velocity, whose product equals $c^{2}$. These velocities are also known as the ‘energy’ and ‘momentum’ velocities, introduced by K. Milton and J. Schwinger. We illustrate general conclusions by explicit calculations for Gaussian and higher-order beams and wavepackets, and also highlight subtleties of the quantum-mechanical description based on the ‘photon wavefunction’.
@@ I Introduction
Subluminal and superluminal velocities in the propagation of light have intrigued scientists for several decades Brillouin (1960). Multiple distinct velocities characterize light propagation, such as phase, group, signal, and energy-transfer velocities. Furthermore, various physical mechanisms can lead to subluminal or superluminal behavior: interactions with matter Chiao and Steinberg (1997); Winful (2006); Boyd and Gauthier (2002); Khurgin (2010); Asano et al. (2016), nontrivial quantum-vacuum effects Latorre, Pascual, and Tarrach (1995), local phase gradients in structured free-space light Berry (2010, 2012); Bliokh et al. (2013), specially engineered space-time wavepackets Bliokh and Nori (2012a); Kondakci and Abouraddy (2019); Yessenov et al. (2022), and the tilt of plane-wave Fourier components in transversely confined light beams Horváth et al. (1996); Porras (2002); Porras, Gonzalo, and Mondello (2003); Giovannini et al. (2015); Bouchard et al. (2016); Gouesbet and Lock (2016); Bareza and Hermosa (2016); Alfano and Nolan (2016); Saari (2018); Petrov (2019); Bliokh (2023a); Gouesbet and Ambrosio (2026).
In this work, we address the simplest scenario among these, namely, the free-space propagation of electromagnetic Gaussian-like wavepackets without engineered internal structure. The only requirement is suitable spatial confinement of the wavepacket, which is necessarily entails a corresponding spread in momentum (wavevector) space. As theoretically described and experimentally observed in Giovannini et al. (2015); Gouesbet and Lock (2016); Alfano and Nolan (2016); Bouchard et al. (2016); Bareza and Hermosa (2016); Saari (2018); Bliokh (2023a), such wavepackets or single photons propagate with a subluminal group velocity $v_{g}<c$, determined by the transverse confinement of the wavepacket. In addition, the wavepacket can also be characterized by a superluminal phase velocity $v_{ph}>c$, such that $v_{g}v_{ph}=c^{2}$, see Fig. 1.
We trace the origin of this result through several complementary approaches and establish useful connections between relativistic field theory, quantum mechanics, and previous studies of this phenomenon. We also examine the extension of the main results to higher-order modes, such as Laguerre-Gaussian beams Bouchard et al. (2016); Bareza and Hermosa (2016).
@@ III Classical wavepacket considerations
This relation is fully analogous to Eq. (4) for the ‘momentum velocity’, so that $\langle{\bf v}_{ph}\rangle\cdot\langle{\bf v}_{g}\rangle=c^{2}$ and $\langle{\bf v}_{ph}\rangle={\bf v}_{P}$.
@@
where $z_{R}=kw_{0}^{2}/2$ is the Rayleigh range, i.e., the characteristic longitudinal scale of beam diffraction. Equation (8) means that a paraxial Gaussian wavepacket accumulates a half-wavelength retardation after propagating a distance $2\pi z_{R}$, i.e., about six Rayleigh ranges, see Fig. 2. This estimate agrees with Refs. Giovannini et al. (2015); Bareza and Hermosa (2016); Gouesbet and Lock (2016), but the retardation effect calculated in Ref. Bliokh (2023a) is twice as small. This difference arises because Ref. Bliokh (2023a) considered a 2D Gaussian wavepacket in the $(x,z)$ plane, whereas here we treat a fully 3D beam. The number of transverse dimensions is doubled, and so is the resulting correction.
@@
$\psi(r_{\perp},z,t)\propto\frac{w_{0}}{w(z)}\exp\!\left(-\frac{r_{\perp}^{2}}{w^{2}(z)}\right)\times$
$\exp\!\left[ikz+i\frac{kr_{\perp}^{2}z}{2(z_{R}^{2}+z^{2})}-i\arctan\!\left(\frac{z}{z_{R}}\right)-i\omega t\right]\!.$ (10)
@@ IV Quantum-mechanical formalism
where the bra-ket inner product implies spatial integration, and we used ${\bf F}^{*}\cdot{\bf F}\equiv|{\bf F}|^{2}=({\bf E}^{2}+{\bf H}^{2})/2=U$. Thus, Eq. (15) is fully equivalent to Eq. (2) for the energy-centroid velocity: $\langle{\bf v}\rangle={\bf v}_{E}$.
@@ V Peculiarities related to the photon wavefunction
VI Higher-order modes
In Section III, we evaluated the average group and phase velocities for the basic Gaussian-like scalar wave modes. One can also consider higher-order modes (e.g., Laguerre-Gaussian or Hermite-Gaussian), which can further enhance the subluminal-group and superluminal-phase effects Bouchard et al. (2016); Bareza and Hermosa (2016).
As an example, consider Laguerre-Gaussian beams Siegman (1986):
$\psi_{LG}\propto\frac{w_{0}}{w}\!\left(\frac{\sqrt{2}r}{w}\right)^{|\ell|}L_{p}^{|\ell|}\!\left(\frac{2r^{2}}{w^{2}}\right)\!\exp\!\left(-\frac{r_{\perp}^{2}}{w^{2}}\right)e^{i\Phi},$ (24)
$\Phi=kz+\frac{kr_{\perp}^{2}z}{2(z_{R}^{2}+z^{2})}-(N+1)\arctan\!\left(\frac{z}{z_{R}}\right)+\ell\varphi-\omega t,$
where $w=w_{0}\sqrt{1+z^{2}/z_{R}^{2}}$, $L_{p}^{|\ell|}$ are generalized Laguerre polynomials, $\ell=0,\pm 1,\pm 2,$... is the azimuthal index, $p=0,1,2,$... is the radial index, $N=|\ell|+2p$, and $\varphi$ is the azimuthal angle with respect to the $z$-axis. The plane-wave spectra of these beams read
$\tilde{\psi}_{LG}\propto\!\left(\frac{w_{0}k_{\perp}}{\sqrt{2}}\right)^{|\ell|}L_{p}^{|\ell|}\!\left(\frac{w_{0}^{2}k_{\perp}^{2}}{2}\right)\!\exp\!\left(-\frac{k_{\perp}^{2}w_{0}^{2}}{4}\right)\!e^{i\ell\phi},$ (25)
where $\phi$ is the azimuthal angle in $\bf k$}-space with respect to the $k_{z}$-axis.
Substituting Eq. (25) into Eq. (6) and using the integral relations from Ref.Phillips and Andrews (1983), we derive, analogously to Eq. (8),
$\langle{v}_{g\,z}\rangle=c\,\frac{\langle k_{z}\rangle}{k}\ \simeq c\left(1-\frac{N+1}{2kz_{R}}\right)\!.$ (26)
Using the phase in Eq. (24), linearized in $z$ near the focal plane ($|z|\ll z_{R}$), we obtain, similarly to Eqs. (11)–(12),
$\langle v_{ph\,z}\rangle\simeq c\left(1+\frac{N+1}{2kz_{R}}\right),$ (27)
Equations (26) and (27) show that the subluminal-group and superluminal-phase effects are enhanced in higher-order Laguerre-Gaussian beams by a factor of $(N+1)$. This agrees with the calculations and measurements in Bareza and Hermosa (2016); Bouchard et al. (2016). This enhancement is directly related to the increased difference $k-\langle k_{z}\rangle\simeq\langle k_{\perp}^{2}\rangle/2k=(N+1)/2z_{R}$, or, equivalently, to the increased Gouy-phase and phase-front curvature terms with $k\langle r_{\perp}^{2}\rangle/2z_{R}^{2}=(N+1)/2z_{R}$ Phillips and Andrews (1983).
Thus, for generic paraxial wavepackets or beams, the deviations of the group and phase velocities from $c$ can be expressed as Giovannini et al. (2015); Gouesbet and Lock (2016); Bareza and Hermosa (2016); Saari (2018)
$1-\frac{\langle v_{g\,z}\rangle}{c}\simeq\frac{\langle v_{ph\,z}\rangle}{c}-1\simeq\frac{\langle k_{\perp}^{2}\rangle}{2k^{2}}$ (28)
As another example, higher-order Bessel beams McGloin and Dholakia (2005) are characterized by a ring-like $\bf k$}-space spectrum with $k_{\perp}={\rm const}$, $k_{z}=\sqrt{k^{2}-k_{\perp}^{2}}={\rm const}$, and the azimuthal phase factor $e^{i\ell\phi}$. For such beams, Eq. (28) shows that the deviations of the group and phase velocities are determined by $\langle k_{\perp}^{2}\rangle=k_{\perp}^{2}$ and are independent of the beam order $\ell$.
VII Conclusions